OpenStudy (shrutipande9):

solve dy/dx=xy/x^2+3y^2

OpenStudy (mimi_x3):

\[\frac{dy}{dx} = \frac{xy}{x^2+3y^2}\] i think this is possible by \( u =y/x)

OpenStudy (mimi_x3):

\( u = y/x\)

OpenStudy (shrutipande9):

okay cool...thnx...dis is what i wanted 2 know:)

OpenStudy (mimi_x3):

thats all lol?

hartnn (hartnn):

mimi to the rescue!

OpenStudy (mimi_x3):

hehe ;)

OpenStudy (shrutipande9):

yep...i just wanted 2 confirm if i m proceeding in d right direction...

OpenStudy (shrutipande9):

!@Mimi_x3 could u plz check if this ans is correct? logx+x^2/6y^2+log(y/x)=c

OpenStudy (shrutipande9):

@Mimi_x3

OpenStudy (mimi_x3):

i got something retarded though :P

OpenStudy (shrutipande9):

Ok ill check...thnx:)

OpenStudy (shrutipande9):

Do des answers match by sum changes??:p

OpenStudy (shrutipande9):

@hartnn @mimi

OpenStudy (unklerhaukus):

dont forget the constant of integration

OpenStudy (mimi_x3):

ye lol, woops

OpenStudy (unklerhaukus):

\[x^2+3y^2=3xy\ln x+Axy\]

OpenStudy (shrutipande9):

ok thnx...:)@UnkleRhaukus

OpenStudy (shrutipande9):

thnx a ton:)

OpenStudy (shrutipande9):

Thnx 2 all @Mimi_x3 @hartnn @UnkleRhaukus

OpenStudy (unklerhaukus):

ive got a minus sign error in my working

OpenStudy (shrutipande9):

ohhh..i need 2 check again den